LABORATORY OBSERVATION OF THE ROTATIONAL SPECTRUM OF A POTENTIAL INTERSTELLAR SUGAR, ERYTHROSE

I. PENA, C. CABEZAS, A. M. DALY, S. MATA, \underline{J. L. ALONSO}, Grupo de Espectroscopía Molecular (GEM), Edificio Quifima, Laboratorios de Espectroscopía y Bioespectroscopía, Parque Científico, Universidad de Valladolid, 47011 Valladolid, Spain.

The rotational spectrum of erythrose has been recorded in the frequency region 6 – 12 GHz using a chirped-pulse Fourier transform microwave spectrometer (CP-FTMW)^{a} combined with a laser ablation (LA) source.^{b} The investigation of rotational spectra of erythrose is of astrophysical and biological relevance. However, no gas-phase data were available on erythrose. It is syrup at room temperature and vaporization using conventional methods leads to decomposition. A non convetional laser ablation method has been successfully used to vaporize erythrose and two cyclic forms have been observed using rotational spectroscopy. α-erythrose has been found to have rotational constants are A = 2586.8998 (21) MHz, B = 2353.0837 (41) MHz and C = 1773.2378 (18) MHz and β-erythrose is characterized with rotational constants A = 3109.2005 (46) MHz, B = 1856.1298 (15) MHz, C = 1616.5557 (18) MHz.

aG. G. Brown, B. C. Dian, K. O. Douglass, S. M. Geyer, S. T. Shipman, B. H. Pate, Rev. Sci. Instrum. 2008, 79, 053103.

bS. Mata, I. Peña, C. Cabezas, J. C. López, J. L. Alonso, J. Mol. Spectrosc. 2012, 280, 91.